please help me solve this problem thanks!
solve for x
2/x^2-6x+8 - 3/x-4 = 7/x-2

Answer by Alan3354(69443)

please help me solve this problem thanks!
solve for x
2/x^2-6x+8 - 3/x-4 = 7/x-2
x^2-6x+8 = (x-4)*(x-2), so that's the DEN
2/DEN - 3(x-2)/DEN = 7(x-4)/DEN
2 - 3x+6 = 7x-28
8-3x = 7x-28
36 = 10x
x = 3.6